The owner of a pet food store wants to mix birdseed that costs $1.25 per pound with sunflower seeds that cost $0.75 per pound to make 50 pounds of a mixture that costs $1.00 per pound. HOw many pounds of each seed should he use?
Answer by checkley77(12844)

1.25X+.75(50-X)=50*1
1.25X+37.5-.75X=50
.50X=50-37.5
.50X=12.5
X=12.5/.5
X=25 POUNDS OF $1.25 BIRDSEED IS USED.
50-25=25 POUNDS OF $.75 SUNFLOWER SEED IS UDED.
PROOF:
1.25*25+.75*25=50
31.25+18.75=50
50=50